OpenStudy (anonymous):

What is the missing side length of an irregular hexagon with side lengths eleven point seven centimeters, nine point four centimeters, twenty point three centimeters, fifteen centimeters and twenty-one point one centimeters, if the perimeter is equal to the perimeter of a regular pentagon with side length twenty-one centimeters? six point five centimeters nine point three centimeters twenty-seven point five centimeters forty-eight point five centimeters

OpenStudy (lucaz):

the perimeter of the regular pentagon is 5 * 21 = 105cm so for the irregular hexagon: 11.7+9.4+20.3+15+21+(x)=105 solvin for x you get: x=105-77.5=27.5cm

OpenStudy (anonymous):

thx

OpenStudy (lucaz):

no problem